Hydrogen atom in its ground is excited by means of monochromatic radiation of wavelength 975 Å. How many different lines are possible in the resulting spectrum ? Calculate the longest wavelength amongst them. You may assume the ionization energy for hydrogen atom as 13.6 eV.

Grade:11

1 Answers

Kevin Nash
9 years ago
Hello Student,
Please find the answer to your question
E = 12400 / λ (in Å) eV = 12400 / 975 = 12.75 eV … (i)
Also
13.6 [ 1 / n21 – 1 / n22] = 12. 75 ⇒ [1 /1 – 1 / n22] = 12.75 / 13. 6 ⇒ n2 = 4
For every possible transition one downward arrow is shown therefore the possibilities are 6.
Note : For longest wavelength, the frequency should be smallest.
This corresponds to the transition from n = 4 to n = 3, the energy will be E4 = 13.6 / 42 ; E3 = - 13.6 / 32
∴ E4 – E3 = 13.6 / 42 – (- 13.6 / 32) = 13.6 [1/9 – 1/16]
= 0.66 eV = 0.66 x 1.6 x 10-19 J = 1.056 x 10-19 J
Now, E = 12400 / λ (inÅ) eV ∴ λ = 18787 Å
Thanks
Kevin Nash
askIITians Faculty